OpenStudy (anonymous):

@abhisar Could you please check my answer? I got 398

OpenStudy (anonymous):

You're correct, great job!

OpenStudy (anonymous):

Thanks

OpenStudy (abhisar):

Yes @lilia222 you are correct. Radius of the inner cylinder = 4-0.75= 3.25 Volume of the cylinder = \(\sf \pi \times r^2 \times h\) Plug in the values and you will get the answer which is 398

OpenStudy (anonymous):

Thank you!

OpenStudy (abhisar):

You're most welcome :)